OpenStudy (el_tucan):

Hi there, here is my question: x = (tan^2)(y^-1)

OpenStudy (anonymous):

y=224

OpenStudy (anonymous):

What to do with it?

OpenStudy (el_tucan):

sorry, differentiate

OpenStudy (anonymous):

in respect to what?

OpenStudy (el_tucan):

in respect to dy/dx, right?

OpenStudy (anonymous):

Well, this is what I'm asking, Taking derivative in respect to x? \(\dfrac{d}{dx}\)?

OpenStudy (el_tucan):

yes

OpenStudy (el_tucan):

I know that y = arctan(x) and x = tan(y) but this is different

OpenStudy (anonymous):

Just making sure. This is implicit differentiation. Just use chain rule then isolate y' (aka \(\dfrac{dy}{dx}\)). You can do it, right?

OpenStudy (el_tucan):

let me try that thanks